At Roccos Restaurant, a large pizza costs $12 plus $1.25 for each additional topping. At pizza palace , a large pizza costs $15 plus $0.75 for each additional topping. For how many toppings will the cost be the same? What will that cost be?


Answer by dabanfield(803)

At Roccos Restaurant, a large pizza costs $12 plus $1.25 for each additional topping. At pizza palace , a large pizza costs $15 plus $0.75 for each additional topping. For how many toppings will the cost be the same? What will that cost be?
Let x be the number of toppings. Then we have:
12 + 1.25*x = 15 + .75*x
1.25*x - .75*x = 15 - 12
.50*x = 3
x = 6